Home Clean room MedAustron HEPHY testbeams old
testbeam-DESY-Oct2017 testbeam-DESY-April2017 testbeam-SPS2015 testbeam-SPS2014 testbeam-DESY14 testbeam-SPS12 testbeam-SPS11 testbeam-SPS10 testbeam-SPS09 testbeam-SPS08 LP-TPC
  Alibava/Mapsa testbeam at DESY in April 2017 (Axel, Grossi), Page 3 of 7  Not logged in ELOG logo
ID Date Author Project Subject Run Number Events XYPosition Cooling Data Text Attachments
  107   Mon Apr 17 17:02:34 2017 Axel KoenigALiBaVa (IFX)run_003_HPK_CB_200V.dat003250000107, 2.6, 230NoUnknownRemarks: Telescope run 606. Angle scan.

Telescope chiller turned off (could
be the root cause of runs crashing).
  
  106   Mon Apr 17 16:36:56 2017 Axel KoenigALiBaVa (IFX)run_002_HPK_CB_200V.dat002450000107, 5.8, 240NoUnknownRemarks: Telescope run 605. Bias scan 200
V. Only one chip.

Telescope chiller turned off (could
  
  105   Mon Apr 17 13:30:01 2017 Axel KoenigALiBaVa (IFX)run_001_HPK_CB_200V.dat001450000107, 0, 240NoUnknownRemarks: Telescope run 604. Bias scan 200
V. Only one chip.

Telescope chiller turned off (could
  
  103   Mon Apr 17 12:26:33 2017 Johannes GrossmannMaPSAHV_Scan45600NoUnknownChanged HV from 60 V to 120 V

12:24 17.4.2017 
  
  104   Mon Apr 17 12:06:11 2017 Axel KoenigALiBaVa (IFX)run_023_Baby_150V.dat023112000102.5, 22.5, 240NoUnknownRemarks: Telescope run 600. Bias scan 150
V.

Telescope chiller turned off (could
  
  101   Mon Apr 17 11:34:55 2017 Axel KoenigALiBaVa (IFX)run_022_Baby_140V.dat022112000102.5, 22.5, 240NoUnknownRemarks: Telescope run 599. Bias scan 140
V.

Telescope chiller turned off (could
  
  100   Mon Apr 17 11:04:40 2017 Axel KoenigALiBaVa (IFX)run_021_Baby_130V.dat021112000102.5, 22.5, 240NoUnknownRemarks: Telescope run 598. Bias scan 130
V.

Telescope chiller turned off (could
  
  99   Mon Apr 17 10:36:22 2017 Axel KoenigALiBaVa (IFX)run_020_Baby_120V.dat020112000102.5, 22.5, 240NoUnknownRemarks: Telescope run 597. Bias scan 120
V.

Telescope chiller turned off (could
  
  98   Mon Apr 17 10:14:15 2017 Axel KoenigcommonRun summary Baby   NoUnknown


Alibava
  
  97   Mon Apr 17 10:10:33 2017 Johannes GrossmannMaPSAHV_Scan38700NoUnknownChanged HV from 80 V to 60 V

Not in condition data applied in
the settings
  
  96   Mon Apr 17 10:05:57 2017 Axel KoenigALiBaVa (IFX)run_019_Baby_110V.dat019112000102.5, 22.5, 240NoUnknownRemarks: Telescope run 596. Bias scan 110
V.

Telescope chiller turned off (could
  
  95   Mon Apr 17 09:29:40 2017 Axel KoenigALiBaVa (IFX)run_018_Baby_90V.dat018112000102.5, 22.5, 240NoUnknownRemarks: Telescope run 595. Bias scan 90
V.

Telescope chiller turned off (could
  
  94   Mon Apr 17 09:04:02 2017 Axel KoenigALiBaVa (IFX)run_006_Baby_100V.dat006k.a.102.5, 22.5, 230NoUnknownRemarks: Telescope run 570. Telescope chiller
turned off (could be the root cause of runs
crashing).
  
  93   Mon Apr 17 09:01:09 2017 Axel KoenigALiBaVa (IFX)run_017_Baby_80V.dat017112000102.5, 22.5, 240NoUnknownRemarks: Telescope run 594. Bias scan 80
V.

Telescope chiller turned off (could
  
  92   Mon Apr 17 08:14:24 2017 Axel KoenigALiBaVa (IFX)run_016_Baby_75V.dat016112000102.5, 22.5, 240NoUnknownRemarks: Telescope run 593. Bias scan 75
V.

Telescope chiller turned off (could
  
  91   Sun Apr 16 22:42:51 2017 Johannes GrossmannMaPSAHV_Scan38700NoUnknownChanged HV from 100V to 80 V

22:41 16.4.2017
  
  90   Sun Apr 16 21:43:15 2017 Axel KoenigALiBaVa (IFX)run_015_Baby_100V.dat0152500000102.5, 22.5, 240NoUnknownRemarks: Telescope run 592. Long run, bias
voltage 100 V. Long run finished without
any problems.
  
  88   Sun Apr 16 20:22:27 2017 Axel KoenigALiBaVa (IFX)run_014_Baby_70V.dat014250000102.5, 22.5, 240NoUnknownRemarks: Telescope run 591. Bias scan 70
V.

Telescope chiller turned off (could
  
  87   Sun Apr 16 19:46:01 2017 Axel KoenigALiBaVa (IFX)run_013_Baby_65V.dat013112000102.5, 22.5, 240NoUnknownRemarks: Telescope run 590. Bias scan 65
V.

Telescope chiller turned off (could
  
  86   Sun Apr 16 19:04:41 2017 Axel KoenigALiBaVa (IFX)run_012_Baby_60V.dat012135000102.5, 22.5, 240NoUnknownRemarks: Telescope run 589. Bias scan 60
V.

Telescope chiller turned off (could
  
ELOG V3.1.4-966e3dd